OpenStudy (anonymous):

I've found out that the limit of the equation below is -5. But I'm not sure how to start when proving that limit with the epsilon-delta definition?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}(-5)\]

OpenStudy (amistre64):

the limit of a constant is the constant ... if you really want, let the equation be -5 + 0x

OpenStudy (amistre64):

find the largest d such that: 0<|x-2|<d given 0< |f(x) - L| < e

OpenStudy (amistre64):

-e < f(x) - L < e L-e < f(x) < L+e convert the range into the domain -5-e < -5+0x < -5+e -e < 0x < e i would say this is true for all values of x but im rusty on my e,d definitions

OpenStudy (amistre64):

what is your f(x)?

OpenStudy (anonymous):

f(x)= -5 ?

OpenStudy (amistre64):

yeah, thats what it looks like from the posting, im jsut not sure if this is starting i the middle of a question where you found the limit as -5 or if its the start of it. it feels like a middle of the line setup

OpenStudy (anonymous):

I'm always used to working with regular equations like 7x+70 for example. But this problem seemed unfamiliar to me.

OpenStudy (amistre64):

graphically, its a sinch  delta = 0 to infinity, an we remain within some epsilon of f(x)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

I want to try to get to absolute value of f(x) - L < e

OpenStudy (amistre64):

so show me how you would do: f(x) = 2x + 1, as x approaches 2

OpenStudy (anonymous):

as "x" approaches 2, the limit would be 5 \[\left| 2x+1 \right|-5<\epsilon \]

OpenStudy (amistre64):

|f(x)-5| not |f(x)| - 5, or am i mistaken?

OpenStudy (anonymous):

\[\left| 2x-4 \right|<\epsilon\]

OpenStudy (anonymous):

\[\left| 2(x-2) \right|<\epsilon\]

OpenStudy (anonymous):

\[\left| 2 \right|*\left| (x-2) \right|<\epsilon/2\]

OpenStudy (amistre64):

|x-2| < e/2 but yeah

OpenStudy (anonymous):

right

OpenStudy (amistre64):

-e/2 < x-2 < e/2 (4-e)/2 < x < (4+e)/2 is valid for all x, and if we know an epsilon, we can easily pick delta for it. right?

OpenStudy (amistre64):

how does the epsilon-delta help us define the limit ... as x to 2?

OpenStudy (amistre64):

does our definition say \[|f(x)-L|<\epsilon\] or can epsilon not equate to zero? its been awhile

OpenStudy (anonymous):

yep. that's it. that's the form I want to get it into to finish verifying the limit.

OpenStudy (amistre64):

http://www.quora.com/What-is-an-epsilon-delta-proof-of-the-Constant-Rule-of-Limits this is as proofy as i can find for it

OpenStudy (amistre64):

f(x) + 5 is 0 which is less than all epsilon greater than 0

OpenStudy (anonymous):

thanks for the link. I will look into that

OpenStudy (amistre64):

other than that, im just not in practice to be able to express it that well.

OpenStudy (amistre64):

good luck tho

OpenStudy (anonymous):

O ok. thanks for at least trying, bud. I'll figure it out